\begin{thebibliography}{10}

\bibitem{kminwise}
Feigenblat, G., Porat, E., Shiftan, A.:
\newblock Exponential time improvement for min-wise based algorithms.
\newblock In: To appear SODA '11: Proceedings of the 22nd annual ACM-SIAM
  symposium on Discrete algorithms, Society for Industrial and Applied
  Mathematics (2011)

\bibitem{patrascu10kwise-lb}
P{\v a}tra{\c s}cu, M., Thorup, M.:
\newblock On the $k$-independence required by linear probing and minwise
  independence.
\newblock In: Proc. 37th International Colloquium on Automata, Languages and
  Programming (ICALP). (2010)  715--726

\bibitem{sizeEstimationFramework}
Cohen, E.:
\newblock Size-estimation framework with applications to transitive closure and
  reachability.
\newblock J. Comput. Syst. Sci. \textbf{55}(3) (1997)  441--453

\bibitem{Datar02estimatingrarity}
Datar, M., Muthukrishnan, S.:
\newblock Estimating rarity and similarity over data stream windows.
\newblock In: In Proceedings of 10th Annual European Symposium on Algorithms,
  volume 2461 of Lecture Notes in Computer Science. (2002)  323--334

\bibitem{NearDuplicate}
Broder, A.Z.:
\newblock Identifying and filtering near-duplicate documents.
\newblock In: COM '00: Proceedings of the 11th Annual Symposium on
  Combinatorial Pattern Matching, London, UK, Springer-Verlag (2000)  1--10

\bibitem{SimilaritySearch}
Haveliwala, T.H., Gionis, A., Klein, D., Indyk, P.:
\newblock Evaluating strategies for similarity search on the web.
\newblock In: WWW '02: Proceedings of the 11th international conference on
  World Wide Web, New York, NY, USA, ACM (2002)  432--442

\bibitem{DBLP:conf/podc/CohenK07}
Cohen, E., Kaplan, H.:
\newblock Summarizing data using bottom-k sketches.
\newblock In: PODC. (2007)  225--234

\bibitem{DBLP:journals/pvldb/CohenK08}
Cohen, E., Kaplan, H.:
\newblock Tighter estimation using bottom k sketches.
\newblock PVLDB \textbf{1}(1) (2008)  213--224

\bibitem{DBLP:conf/spire/BachrachHP09}
Bachrach, Y., Herbrich, R., Porat, E.:
\newblock Sketching algorithms for approximating rank correlations in
  collaborative filtering systems.
\newblock In Karlgren, J., Tarhio, J., Hyyr{\"o}, H., eds.: SPIRE. Volume 5721
  of Lecture Notes in Computer Science., Springer (2009)  344--352

\bibitem{Bachrach:2009:IJCAI}
Bachrach, Y., Porat, E., Rosenschein, J.S.:
\newblock Sketching techniques for collaborative filtering.
\newblock In: The Twenty-First International Joint Conference on Artificial
  Intelligence (IJCAI 2009), Pasadena, California (2009)  2016--2021

\bibitem{1242592}
Manku, G.S., Jain, A., Das~Sarma, A.:
\newblock Detecting near-duplicates for web crawling.
\newblock In: WWW '07: Proceedings of the 16th international conference on
  World Wide Web, New York, NY, USA, ACM (2007)  141--150

\bibitem{1115528}
Cormode, G., Muthukrishnan, S.:
\newblock What's new: finding significant differences in network data streams.
\newblock IEEE/ACM Trans. Netw. \textbf{13}(6) (2005)  1219--1232

\bibitem{872790}
Ganguly, S., Garofalakis, M., Rastogi, R.:
\newblock Processing set expressions over continuous update streams.
\newblock In: SIGMOD '03: Proceedings of the 2003 ACM SIGMOD international
  conference on Management of data, New York, NY, USA, ACM (2003)  265--276

\bibitem{1242610}
Das, A.S., Datar, M., Garg, A., Rajaram, S.:
\newblock Google news personalization: scalable online collaborative filtering.
\newblock In: WWW '07: Proceedings of the 16th international conference on
  World Wide Web, New York, NY, USA, ACM (2007)  271--280

\bibitem{378687}
Gibbons, P.B., Tirthapura, S.:
\newblock Estimating simple functions on the union of data streams.
\newblock In: SPAA '01: Proceedings of the thirteenth annual ACM symposium on
  Parallel algorithms and architectures, New York, NY, USA, ACM (2001)
  281--291

\bibitem{511502}
Haveliwala, T.H., Gionis, A., Klein, D., Indyk, P.:
\newblock Evaluating strategies for similarity search on the web.
\newblock In: WWW '02: Proceedings of the 11th international conference on
  World Wide Web, New York, NY, USA, ACM (2002)  432--442

\bibitem{1148243}
Yang, H., Callan, J.:
\newblock Near-duplicate detection by instance-level constrained clustering.
\newblock In: SIGIR '06: Proceedings of the 29th annual international ACM SIGIR
  conference on Research and development in information retrieval, New York,
  NY, USA, ACM (2006)  421--428

\bibitem{1148222}
Henzinger, M.:
\newblock Finding near-duplicate web pages: a large-scale evaluation of
  algorithms.
\newblock In: SIGIR '06: Proceedings of the 29th annual international ACM SIGIR
  conference on Research and development in information retrieval, New York,
  NY, USA, ACM (2006)  284--291

\bibitem{Mulmuley92}
Mulmuley, K.:
\newblock Randomized geometric algorithms and pseudo-random generators.
\newblock In: SFCS '92: Proceedings of the 33rd Annual Symposium on Foundations
  of Computer Science, Washington, DC, USA, IEEE Computer Society (1992)
  90--100

\bibitem{MinwiseIndependentPermutations}
Broder, A.Z., Charikar, M., Frieze, A.M., Mitzenmacher, M.:
\newblock Min-wise independent permutations (extended abstract).
\newblock In: STOC '98: Proceedings of the thirtieth annual ACM symposium on
  Theory of computing, New York, NY, USA, ACM (1998)  327--336

\bibitem{Broder97onthe}
Broder, A.Z.:
\newblock On the resemblance and containment of documents.
\newblock In: In Compression and Complexity of Sequences (SEQUENCES'97, IEEE
  Computer Society (1997)  21--29

\bibitem{SyntacticClustering}
Broder, A.Z., Glassman, S.C., Manasse, M.S., Zweig, G.:
\newblock Syntactic clustering of the web.
\newblock In: Selected papers from the sixth international conference on World
  Wide Web, Essex, UK, Elsevier Science Publishers Ltd. (1997)  1157--1166

\bibitem{Indyk99asmall}
Indyk, P.:
\newblock A small approximately min-wise independent family of hash functions.
\newblock In: Journal of Algorithms. (1999)  454--456

\bibitem{Saks99lowdiscrepancy}
Saks, M., Srinivasan, A., Zhou, S., Zuckerman, D.:
\newblock Low discrepancy sets yield approximate min-wise independent
  permutation families.
\newblock In: In Proc. International Workshop on Randomization and
  Approximation Techniques in Computer Science, Springer (1999)  29--32

\end{thebibliography}

\documentclass[envcountreset]{llncs}
\usepackage{amsmath}    
\usepackage{multirow}
\newcommand{\ekmin}{\emph{$\epsilon$-k-min-wise }}
\newcommand{\edkmin}{\emph{$\epsilon$-d-k-min-wise }}
\newcommand{\kmin}{\emph{k-min-wise }}
\newcommand{\dkmin}{\emph{d-k-min-wise }}
\newcommand{\minwise}{\emph{min-wise }}
\newcommand{\eminwise}{\emph{$\epsilon$-min-wise }}

\numberwithin{lemma}{section}
\numberwithin{theorem}{section}
\numberwithin{table}{section}
\numberwithin{definition}{section}

\begin{document}

\title{Even Better Framework for \minwise Based Algorithms}

\author{Guy Feigenblat, Ely Porat and Ariel Shiftan}

\institute{Department of Computer Science, Bar-Ilan University, Ramat Gan 52900, Israel \\ \email{\{feigeng, porately, shiftaa\}@cs.biu.ac.il}
}

\maketitle


\begin{abstract}
In a recent paper from SODA11 \cite{kminwise} the authors introduced a general framework for exponential time improvement of \minwise based algorithms by defining and constructing almost \kmin independent family of hash functions. Here we take it a step forward and reduce the space and the independent needed for representing the functions, by defining and constructing a \dkmin independent family of hash functions. Surprisingly, for most cases only $8$-wise independent is needed for exponential time and space improvement.
Moreover, we bypass the $O(\log{\frac{1}{\epsilon}})$ independent lower bound for approximately \minwise functions \cite{patrascu10kwise-lb}, as we use alternative definition.
In addition, as the independent's degree is a small constant it can be implemented efficiently.

Informally, under this definition, all subsets of size $d$ of any fixed set $X$ have an equal probability to have hash values among the minimal $k$ values in $X$,
where the probability is over the random choice of hash function from the family.
This property measures the randomness of the family, as choosing a truly random function, obviously, satisfies the definition for $d=k=|X|$.
We define and give an efficient time and space construction of approximately \dkmin independent family of hash functions.
The degree of independent required is optimal, i.e. only $O( d )$ for $2 \le d < k=O(\frac{d}{\epsilon^2})$, where $\epsilon \in (0,1)$ is the desired error bound.
This construction can be used to improve many \minwise based algorithms,
such as \cite{sizeEstimationFramework,Datar02estimatingrarity,NearDuplicate,SimilaritySearch,DBLP:conf/podc/CohenK07,DBLP:journals/pvldb/CohenK08,DBLP:conf/spire/BachrachHP09,Bachrach:2009:IJCAI,1242592,1115528,872790,1242610,378687,511502,1148243,1148222}, as will be discussed here.
To our knowledge such definitions, for hash functions, were never studied and no construction was given before.
\end{abstract}

\section{Introduction}
Hash functions are fundamental building blocks of many algorithms. They map values from one domain to another, usually smaller. Although they have been studied for many years, designing hash functions is still a hot topic in modern research. In a perfect world we could use a truly random hash function, one that would be chosen randomly out of all the possible mappings.

Specifically, consider the domain of all hash functions $h: N \rightarrow M $, where $|N|=n$ and $|M|=m$. As we need to map each of the $n$ elements in the source into one of the $m$ possible mappings, the number of bits needed to maintain each function is $n \log m$. Since nowadays we often have massive amount of data to process, this amount of space is not feasible. Nevertheless, most algorithms do not really need such a high level of randomness, and can perform well enough with some relaxations. In such cases one can use a much smaller domain of hash functions. A smaller domain implies lesser space requirement at the price of a lower level of randomness.

As an illustrative example, the notion of \emph{2-wise-independent} family of hash functions assures the independence of each pair of elements. It is known that only $2\log m$ bits are enough in order to choose and maintain such a function out of the family.

This work is focused on the area of min-hashing. One derivative of min-hashing is \minwise independent permutations, which were first introduced in \cite{Mulmuley92,MinwiseIndependentPermutations}. A family of {\bf permutations} $F \in S_{n} $ (where $S_{n}$ the symmetric group) is {\bf min-wise independent} if for any set $ X \subseteq [n]$ (where $[n]=\{0,\dots,n-1\} $) and any $x \in X$, where $\pi $ is chosen uniformly at random in $F$, we have:
\begin{equation}
 Pr  [\min \{\pi(X)\} = \pi(x)] = \frac{1}{|X|}  \nonumber
\end{equation}

Similarly, a family of {\bf functions} $\mathcal{H} \in [n] \rightarrow [n] $ (where $[n]=\{0,\dots,n-1\} $) is called {\bf min-wise independent} if for any $ X \subseteq [n]$, and for any $x \in X$, where $h$ is chosen uniformly at random in $\mathcal{H}$, we have:
\begin{equation}
 Pr _{h \in \mathcal{H}} [\min \{h(X)\} = h(x)] = \frac{1}{|X|}  \nonumber
\end{equation}

Min hashing is a widely used tool for solving problems in computer science such as estimating similarity \cite{MinwiseIndependentPermutations,Broder97onthe,SyntacticClustering}, rarity \cite{Datar02estimatingrarity},  transitive closure \cite{sizeEstimationFramework}, web page duplicate detection \cite{NearDuplicate,1242592,1148243,1148222}, sketching techniques \cite{DBLP:journals/pvldb/CohenK08,DBLP:conf/podc/CohenK07}, and other data mining problems \cite{SimilaritySearch,1242610,511502,DBLP:conf/spire/BachrachHP09,Bachrach:2009:IJCAI}.

One of the key properties of min hashing is that it enables to sample the universe of the elements being hashed.
This is because each element, over the random choice of hash function out of the family, has equal probability of being mapped to the minimal value, regardless of the number of occurrences of the element. Thus, by maintaining the element with the minimal hash value over the input, one can sample the universe.

Similarity estimation of data sets is a fundamental tool in mining data. It is often calculated using the Jaccard similarity coefficient which is defined by $ \frac{ |A \cap B|}{|A \cup B|} $, where $A$ and $B$ are two data sets.
By maintaining the minimal hash value over two sets of data inputs $A$ and $B$, the probability of getting the same hash value is exactly $ \frac{ |A \cap B|}{|A \cup B|} $, which equals the Jaccard similarity coefficient, as described in \cite{MinwiseIndependentPermutations,Broder97onthe,SyntacticClustering,sizeEstimationFramework}. \\

Indyk in \cite{Indyk99asmall} was first to give a construction of a small approximately \minwise independent family of hash functions, another construction was proposed in \cite{Saks99lowdiscrepancy}. \\
A family of functions $ \mathcal{H} \subseteq [n] \rightarrow [n] $ is called {\bf approximately min-wise independent}, or \eminwise independent, if,  for any $X \subseteq  [n]$, and for any $x \in X$, where $h$ is chosen uniformly at random in $\mathcal{H}$, we have:
 \begin{equation}
 Pr _{h \in \mathcal{H}} [\min \{h(X)\} = h(x)] = \frac{1}{|X|} (1 \pm \epsilon) \nonumber
\end{equation}
where $\epsilon \in (0,1)$ is the desired error bound, and $O( \log (\frac{1}{\epsilon}) )$ independent is needed.
P{\v a}tra{\c s}cu and Thorup showed in \cite{patrascu10kwise-lb} that $O(\log{\frac{1}{\epsilon}})$ independent is needed for maintaining an approximately \minwise function, hence Indyk's construction is optimal,
and the minimal number of bits needed to represent each function is $O( \log n \log (\frac{1}{\epsilon}) )$. \\

In a previous paper \cite{kminwise} the authors defined and gave a construction for
{\bf approximately \emph{\kmin (\ekmin)} independent} family of hash functions: \\
A family of functions $\mathcal{H} \subseteq [n]\rightarrow [n]$ (where $[n] = \{ 0 \ldots n-1\}$) is called \emph{\ekmin}independent if for any $X\subseteq [n]$ and for any $Y\subset X$, $|Y| = k$ we have
\[
\Pr_{h\in \mathcal{H}}\left[ \max_{y \in Y} h(y) < \min_{z \in X-Y} h(z)\right] = \frac{1}{ {|X| \choose |Y|}}(1\pm \epsilon),
\]
where the function $h$ is chosen uniformly at random from $\mathcal{H}$, and $\epsilon \in(0,1)$ is the error bound. \\
It was also shown in \cite{kminwise} that choosing uniformly at random from $O( k \log\log\frac{1}{\epsilon} +  \log \frac{1}{\epsilon}  )$-wise independent family of hash functions is approximately \kmin independent.
In most applications, $k$ different approximately \minwise hash functions were used, and they proposed to replace them with only one approximately \kmin hash function.
As the $k$ elements are fully independent, the precision is preserved.
Furthermore, this reduces exponentially the running time and asymptotically the space of previous known results for \minwise based algorithms.

\subsection{Our Contribution}
In this paper we define and construct a small approximately \dkmin independent family of hash functions.
First, we extend the notion of \minwise independent family of hash functions by defining a \dkmin independent family of hash functions.
Then, we show a construction of an approximated such family.
Under this definition, all subsets of size $d$ of any fixed set $X$ have an equal probability to have hash values among the minimal $k$ values in $X$,
where the probability is over the random choice of hash function from the family. The formal definition is given in section 2.
The degree of independent and the space needed by our construction is $O(d)$,
for $2 \le d < k=O(\frac{d}{\epsilon^2})$, where $\epsilon \in (0,1)$ is the error bound. 
The construction is optimal, since by our definition the $k$ elements are $d$-wise independent.
In addition, the dependency on $d$ but not on $k$ is surprising, but the intuition behind that is the stability property of the $k$'th sized element, for large enough $k$.
Hence, the randomness needed is mainly for the independence of the $d$ elements.

We argue that for most applications it is sufficient to use constant $d=2$. This yields the need of only $8$-wise independent hash functions, which can be implemented efficiently. Our innovative approach bypasses the $O(\log{\frac{1}{\epsilon}})$ lower bound of approximately \minwise functions, as this constraint is not required by our definition.

To utilize our construction we propose a simple and even better general framework for exponential time improvement of \minwise based algorithms,
such as in \cite{sizeEstimationFramework,Datar02estimatingrarity,NearDuplicate,SimilaritySearch,DBLP:conf/podc/CohenK07,DBLP:journals/pvldb/CohenK08,DBLP:conf/spire/BachrachHP09,Bachrach:2009:IJCAI,1242592,1115528,872790,1242610,378687,511502,1148243,1148222}.
The common between these algorithms is that they use $c>1$ approximately \minwise hash functions in order to sample $c$ elements independently from the universe.
We propose to replace them with fewer (say $O(\log\frac{1}{\tau})$, for $\tau\in(0,1)$) approximately \dkmin (for $k<c$) independent functions, where each samples $k$ elements.
The $k$ elements are $d$-wise independent, therefore we can use Chebyshev's inequality to bound the precision. Specifically, pair-wise is sufficient for applying Chebyshev, and this is why $d=2$ is usually used.
By combining the $\log\frac{1}{\tau}$ samples using Chernoff bound, the precision becomes as desired.
The above procedure does not change the algorithm itself, but only the way it samples, hence it is simple to adapt.
We found this to improve exponentially the time and space complexity (as the space for each function is constant).\\





\subsection{Outline}
The outline of the paper is as follows.
In section 2 we define the notion of \dkmin and approximately \dkmin independent families. 
In section 3 we present a construction of such family.
Finally, few lemmata are left to the appendix.

\section{Definitions}
We will start by giving our definitions for exact and approximately \dkmin independent family of hash functions, which are generalization of \minwise independent family of hash functions.
\begin{definition}
For any set $X$ we define $MIN_k(X)$ to be the set of $k$ smallest elements in $X$.
\end{definition}
\begin{definition}
For any set $X$ we define $RANK_k(X)$ to be the $k$-th elements in $X$, where the elements are sorted by value.
\end{definition}
\begin{definition}
For any set $X$, and hash function $h$ we define $h(X)$ to be the set of all hash values of all elements in $X$.
\end{definition}

\begin{definition}
\emph{\dkmin}independent family of hash functions: \\
A family of functions $\mathcal{H} \subseteq [u]\rightarrow [u]$ (where $[u] = \{ 0 \ldots u-1\}$) is called \emph{\dkmin}independent if for any $X\subseteq [u]$, $|X|=n-d$, for any $Y\subseteq [u]$, $|Y| = d$, $X\bigcap Y=\emptyset$, $d<=k$, we have
\begin{equation}
\Pr_{h\in \mathcal{H}}\left[ RANK_d(h(Y)) <  RANK_{k-d+1}(h(X))\right] = \frac{{k \choose d}}{ {n \choose d}}  \nonumber
\end{equation}
Where the function $h$ is chosen uniformly at random from $\mathcal{H}$.
\end{definition}

\begin{definition}
Approximately \emph{\dkmin (\edkmin)} independent family of hash functions: \\
A family of functions $\mathcal{H} \subseteq [u]\rightarrow [u]$ (where $[u] = \{ 0 \ldots u-1\}$) is called approximately \emph{\dkmin}independent if for any $X\subseteq [u]$, $|X|=n-d$, for any $Y\subseteq [u]$, $|Y| = d$, $X\bigcap Y=\emptyset$, $d<=k$, we have
\begin{equation}
\Pr_{h\in \mathcal{H}}\left[ RANK_d(h(Y)) <  RANK_{k-d+1}(h(X))\right] = \frac{{k \choose d}}{ {n \choose d}} (1\pm \epsilon) \nonumber
\end{equation}
Where the function $h$ is chosen uniformly at random from $\mathcal{H}$, and $\epsilon \in(0,1)$ is the error bound.
\label{def2}
\end{definition}

\section{Construction}
In this section we provide a construction for approximately \dkmin independent family of hash functions.
Let $\Pr\left[\cdot\right]$ denote a fully random probability measure over $[u]\rightarrow [u]$, and let $\Pr_l\left[\cdot\right]$ denote any l-wise independent probability measure over the same domain. We divide the universe into $|\phi|$ non-overlapping blocks, which will be defined in the next section.

\begin{lemma}
\label{lemma1}
\begin{equation}
\label{a1}
\Pr \left[ h(y_1), h(y_2),\dots,h(y_d)  < RANK_{k-d+1}(h(X))\right] =  \frac{k}{n}\frac{k-1}{n-1}\dots\frac{k-d+1}{n-d+1} \nonumber
\end{equation}
\end{lemma}
\begin{proof}
Consider $n$ ordered elements divided into two groups --- one of size $n-d$, and the other of size $d$. The number of possible locations of the $d$ elements is ${n \choose d}$. There are ${k \choose d}$ possible locations in which the $d$ elements are among the $k$ smallest elements. Hence, the probability for the $d$ element to be among the $k$'th smallest elements is:
\begin{equation}
\label{a1}
\Pr \left[ h(y_1), h(y_2),\dots,h(y_d)  < RANK_{k-d+1}(h(X))\right] =
\frac{{k \choose d}}{{n \choose d}} =  \frac{k}{n}\frac{k-1}{n-1}\dots\frac{k-d+1}{n-d+1} \nonumber
\end{equation}
\qed
\end{proof}

Since the blocks in $\phi$ are non-overlapping \\ $\sum_{i\in\phi} \Pr_l\left[RANK_{k-d+1}(h(X)) \in b_i\right]=1$, using lemma \ref{lemma1} we get
\begin{equation}
\label{a1}
\Pr \left[ h(y_1), h(y_2),\dots,h(y_d)  < RANK_{k-d+1}(h(X))\right] = \nonumber
\end{equation}
\begin{equation}
\frac{k}{n}\frac{k-1}{n-1}\dots\frac{k-d+1}{n-d+1} \sum_{i\in\phi} \Pr_l\left[RANK_{k-d+1}(h(X)) \in b_i\right] \nonumber
\end{equation}

\begin{lemma}
\label{lemma2}
Let $\Delta=$
\begin{equation}
\sum_{i\in\phi} \Pr_l\left[ RANK_{k-d+1}(h(X)) \in b_i\right] \times \nonumber
\end{equation}
\begin{equation}
\left[ \Pr_l\left[ h(y_1), \dots,h(y_d) \le RANK_{k-d+1}(h(X)) \ | \ RANK_{k-d+1}(h(X)) \in b_i  \right] - \frac{k}{n}\frac{k-1}{n-1}\dots\frac{k-d+1}{n-d+1} \right] \nonumber
\end{equation}
Any family of $l$-wise independent is approximately \emph{\dkmin}independent if
\begin{equation}
 -\epsilon\frac{{k \choose d}}{{n \choose d}} \le \Delta \le \epsilon\frac{{k \choose d}}{{n \choose d}} \nonumber
\end{equation}
\end{lemma}
\begin{proof}
Based on the complete probability formula, in the l-wise independent case
\begin{equation}
\label{a2}
\Pr_l \left[ h(y_1), h(y_2),\dots,h(y_d) < RANK_{k-d+1}(h(X))\right] = \nonumber
\end{equation}
\begin{equation}
\sum_{i\in\phi} \Pr_l\left[ RANK_{k-d+1}(h(X)) \in b_i\right]  \cdot \nonumber
\end{equation}
\begin{equation}
 \Pr_l\left[ h(y_1), h(y_2),\dots,h(y_d) < RANK_{k-d+1}(h(X)) \ | \ RANK_{k-d+1}(h(X)) \in b_i  \right] \nonumber
\end{equation}
By definition, any family of $l$-wise independent is approximately \dkmin independent if
\begin{equation}
\Pr_l \left[ h(y_1), h(y_2),\dots,h(y_d) < RANK_{k-d+1}(h(X))\right] = \frac{{k \choose d}}{ {n \choose d}} (1\pm \epsilon) \nonumber
\end{equation}
which is satisfied if
\begin{equation}
 -\epsilon\frac{{k \choose d}}{{n \choose d}} \le \Delta \le \epsilon\frac{{k \choose d}}{{n \choose d}} \nonumber
\end{equation}
\qed
\end{proof}

\subsection{Blocks partitioning}
We divide the universe $\left[0,|U|\right]$ into non-overlapping blocks. Let $t=k-d+1$ and $m = n-d = |X|$, we construct the blocks around $\frac{t|U|}{m}$, as follows:
$b_i = \left[ (1+\epsilon (i-1)) \frac{t|U|}{m}, (1+\epsilon i) \frac{t|U|}{m} \right)$ \\
We refer to blocks $b_i$ for $i > 0$ as 'positive blocks' and 'negative blocks' otherwise ($i \le 0$).
For the rest of the paper, we ignore blocks which are outside $\left[0,|U|\right]$.

\subsection{Bounding $\Pr_l\left[ RANK_{t}(h(X)) \in b_{i}\right]$}

\begin{lemma}
\label{lemma3}
For $i>0$, $d>0$, $\epsilon \in (0,1)$, $k>d-1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2}$ and $l\ge 2d+2$
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_i\right] \le \frac{1}{i^{d+1}} \nonumber
\end{equation}

\end{lemma}
\begin{proof}
For block $b_i$, $X = \{x_1,\dots,x_{m}\} $ we define $Z_j$ to be indicator variable s.t.
\begin{equation}
Z_j =
\left\{
	\begin{array}{ll}
		1 \ \ \ &  h(x_j) < (1+\epsilon j) \frac{t|U|}{m}\\
		0 \ \ \ & otherwise
	\end{array}
\right. \nonumber
\end{equation}
In addition we define $Z = \sum_jZ_j$, and $E_i$ to be the expected value of $Z$. Notice that since $Z$ is sum of indicator variables $E_i= (1+\epsilon i) \frac{t}{m}(m) = t(1+\epsilon i)$. \\
We use the above definitions to show that
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_i\right] \le \nonumber
\end{equation}
\begin{equation}
\Pr_l[ \text{number of hash values smaller than the lower boundary of block } b_i < t] = \nonumber
\end{equation}
\begin{equation}
\Pr_l[Z < t] \le \Pr_l[E_i-Z \ge E_i-t] \le \Pr_l[|E_i-Z| \ge E_i-t] =\nonumber
\end{equation}
\begin{equation}
\Pr_l[|Z-E_i| \ge t(1+\epsilon i)-t] \nonumber
\end{equation}
Using Markov inequality, assuming l is even:
\begin{equation}
\Pr_l[|Z-E_i| \ge t\epsilon i] \le \frac{E(|Z-E_i|^l)}{[t\epsilon i]^l} \nonumber
\end{equation}

We use the following from lemma \ref{lemmaIndyk}:
\begin{equation}
E(|Z-E_i|^l) \le 8(6l)^{\frac{l+1}{2}}(E_i)^{\frac{l}{2}}\nonumber
\end{equation}
Thus,
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_i\right] \le
 \frac{8(6l)^{\frac{l+1}{2}}(t(1+\epsilon i))^{\frac{l}{2}}}{[t\epsilon i]^l} \nonumber
\end{equation}
In order to have
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_i\right] \le \frac{1}{i^{d+1}} \nonumber
\end{equation}
We need
\begin{equation}
\frac{8(6l)^{\frac{l+1}{2}}(t(1+\epsilon i) )^{\frac{l}{2}}}{[t\epsilon i]^l} \le \frac{1}{i^{d+1}} \nonumber
\end{equation}
or
\begin{equation}
\frac{8(6l)^{\frac{l+1}{2}}(1+\epsilon i)^{\frac{l}{2}}}{\epsilon^l i^{l-d-1}} \le t^{\frac{l}{2}} \nonumber
\end{equation}
\begin{equation}
\frac{8^{\frac{2}{l}}(6l)^{1+\frac{1}{l}}(1+\epsilon i)}{\epsilon^2 i^{\frac{2(l-d-1)}{l}}} \le t \nonumber
\end{equation}
choosing $l\ge 2d+2$ and substituting variables, we now need to show
\begin{equation}
\frac{8^{\frac{2}{l}}(6l)^{1+\frac{1}{l}}(1+\epsilon i)}{\epsilon^2 i} + d - 1 \le k \nonumber
\end{equation}
\begin{equation}
8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2 i} + 8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}\mathbf{}\epsilon}{\epsilon^2}+ d - 1\le k \nonumber
\end{equation}
which is satisfied for $k>d - 1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2}$
\qed
\end{proof}

As the proof for negative blocks is similar, we left it to the appendix. See lemma \ref{lemma4}, as follows:\\
\begin{emph}
{\bf Lemma}
For $i\ge 0$, $d>0$, $\epsilon \in (0,1)$, $k> d - 1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2}$ and $l\ge 2d+2$:
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_{-i}\right] \le \frac{1}{i^{d+1}} \nonumber
\end{equation}
\end{emph}

\subsection{Bounding $\Delta$}
In this section we prove the upper and lower bounds of lemma \ref{lemma2}, i.e. that  $-\epsilon\frac{{k \choose d}}{{n \choose d}} \le \Delta \le \epsilon\frac{{k \choose d}}{{n \choose d}}$.
As $d=2$ is the main use case, and since the proof for general $d$ is too technical, we focus on the first, and leave the extended proof for the full paper.

We define $\Pr_l\left[ RANK_{t}(h(X)) \in b_{\ge j}\right]$ by $\Pr_l\left[ \cup_{i=j}^{\infty} RANK_{t}(h(X)) \in b_{i} \right]$ and similarly $\Pr_l\left[ RANK_{t}(h(X)) \in b_{\le j}\right]$ by $\Pr_l\left[ \cup_{i=-\infty}^{j} RANK_{t}(h(X)) \in b_{i} \right]$.

\begin{lemma}
\label{lemmaDeltaUpper}
For $d=2$, $d>0$, $\epsilon \in (0,1)$, $k>d-1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2}$ and $l\ge 2d+2$:\\
\begin{center}
$\Delta \le \epsilon\frac{{k \choose d}}{{n \choose d}}$
\end{center}
\end{lemma}
\begin{proof}

\begin{equation}
\sum_{i=-\infty}^{\infty} \Pr_{l+d}\left[ RANK_{t}(h(X)) \in b_i\right] \times \nonumber
\end{equation}
\begin{equation}
\left[ \Pr_{l+d}\left[ h(y_1), h(y_2),\dots,h(y_d) \le RANK_{t}(h(X)) \ | \ RANK_{t}(h(X)) \in b_i  \right] - \frac{{k \choose d}}{{n \choose d}}\right] \le \nonumber
\end{equation}
Using $d$ independent (out of $l + d$) for $h(y_1), h(y_2),\dots,h(y_d)$, we get
\begin{equation}
\sum_{i=-\infty}^{\infty} \Pr_l\left[ RANK_{t}(h(X)) \in b_i\right] \left[ (\frac{t}{m})^d(1+\epsilon i)^d-\frac{{k \choose d}}{{n \choose d}} \right] = \nonumber
\end{equation}

\begin{equation}
\sum_{i=-\infty}^{-1} \Pr_l\left[ RANK_{t}(h(X)) \in b_i\right] \left[ (\frac{t}{m})^d(1+\epsilon i)^d- \frac{{k \choose d}}{{n \choose d}}\right] + \nonumber
\end{equation}
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_0\right] \left[ (\frac{t}{m})^d-\frac{{k \choose d}}{{n \choose d}} \right] + \nonumber
\end{equation}
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_1\right] \left[ (\frac{t}{m})^d(1+\epsilon)^d-\frac{{k \choose d}}{{n \choose d}} \right] + \nonumber
\end{equation}
\begin{equation}
\sum_{i=2}^{\infty} \Pr_l\left[ RANK_{t}(h(X)) \in b_i\right] \left[ (\frac{t}{m})^d(1+\epsilon i)^d-\frac{{k \choose d}}{{n \choose d}} \right] = \nonumber
\end{equation}
By changing the order we get a telescoping sum as follows:
\begin{equation}
\sum_{i=-\infty}^{-1} \Pr_l\left[ RANK_{t}(h(X)) \in b_{\le i}\right] \left[ (\frac{t}{m})^d(1+\epsilon i)^d- (\frac{t}{m})^d(1+\epsilon (i+1))^d\right] + \nonumber
\end{equation}
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_{\le 0}\right] \left[ (\frac{t}{m})^d-\frac{{k \choose d}}{{n \choose d}} \right] + \nonumber
\end{equation}
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_{\ge 1}\right] \left[ (\frac{t}{m})^d(1+\epsilon)^d-\frac{{k \choose d}}{{n \choose d}} \right] + \nonumber
\end{equation}
\begin{equation}
\sum_{i=2}^{\infty} \Pr_l\left[ RANK_{t}(h(X)) \in b_{\ge i}\right] \left[ (\frac{t}{m})^d(1+\epsilon i)^d- (\frac{t}{m})^d(1+\epsilon (i-1))^d\right] \le \nonumber
\end{equation}

Applying lemma \ref{lemma3} and lemma \ref{lemma4}, bounding the probabilities of blocks $b_0$ and $b_1$ with $1$
\begin{equation}
\sum_{i=-\infty}^{-1} \frac{1}{|i|^{d+1}} |(\frac{t}{m})^d(1+\epsilon i)^d- (\frac{t}{m})^d(1+\epsilon (i+1))^d| + \nonumber
\end{equation}
\begin{equation}
|(\frac{t}{m})^d-\frac{{k \choose d}}{{n \choose d}}| + |(\frac{t}{m})^d(1+\epsilon)^d-\frac{{k \choose d}}{{n \choose d}}| +\nonumber
\end{equation}
\begin{equation}
\sum_{i=2}^{\infty} \frac{1}{i^{d+1}} |(\frac{t}{m})^d(1+\epsilon i)^d- (\frac{t}{m})^d(1+\epsilon (i-1))^d| = \nonumber
\end{equation}

We now substitute $d$

\begin{equation}
\sum_{i=1}^{\infty} \frac{1}{|i|^3} |(\frac{k-1}{n-2})^2(\epsilon^2(2i-1) - 2\epsilon))| + \nonumber
\end{equation}
\begin{equation}
|(\frac{k-1}{n-2})^2-\frac{k}{n}\frac{k-1}{n-1}| + |(\frac{k-1}{n-2})^2(1+\epsilon)^2-\frac{k}{n}\frac{k-1}{n-1}| +\nonumber
\end{equation}
\begin{equation}
\sum_{i=2}^{\infty} \frac{1}{i^3} |(\frac{k-1}{n-2})^2(\epsilon^2(2i-1) + 2\epsilon))| \le \nonumber
\end{equation}

\begin{equation}
2\frac{k}{n}\frac{k-1}{n-1} \sum_{i=1}^{\infty} \frac{1}{|i|^3} |\epsilon^2(2i-1) - 2\epsilon| + \nonumber
\end{equation}
\begin{equation}
|(\frac{k-1}{n-2})^2-\frac{k}{n}\frac{k-1}{n-1}| + |(\frac{k-1}{n-2})^2(1+\epsilon)^2-\frac{k}{n}\frac{k-1}{n-1}| +\nonumber
\end{equation}
\begin{equation}
2\frac{k}{n}\frac{k-1}{n-1} \sum_{i=2}^{\infty} \frac{1}{i^3} |\epsilon^2(2i-1) + 2\epsilon| \le \nonumber
\end{equation}

\begin{equation}
c\frac{k}{n}\frac{k-1}{n-1}\epsilon  \nonumber
\end{equation}

\end{proof}
\qed

The proof for negative blocks is similar, hence we left it to the appendix. See lemma \ref{lemmaDeltaLower}, as follows:\\
\begin{emph}
{\bf Lemma}
For $d=2$, $\epsilon \in (0,1)$, $k>d-1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2}$ and $l\ge 2d+2$: \ \ \ $-\Delta \le \epsilon\frac{{k \choose d}}{{n \choose d}}$.
\end{emph}

\begin{theorem}
For $d=2$, $\epsilon \in (0,1)$, $\epsilon'=\frac{\epsilon}{c}$ , $k>d-1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon'^2}$ and $l\ge 3d+2$,
Any l-wise independent family of hash functions is approximately \emph{\dkmin (\edkmin)}.
\end{theorem}
\begin{proof}
Applying lemma \ref{lemmaDeltaUpper} and lemma \ref{lemmaDeltaLower} to lemma \ref{lemma2} concludes the proof.
\end{proof}
\qed

\begin{theorem}
For $d\le\frac{k}{c'}$, $\epsilon \in (0,1)$, $\epsilon'=\frac{\epsilon}{c}$ , $k>d-1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon'^2}$ and $l\ge 3d+2$,
Any l-wise independent family of hash functions is approximately \emph{\dkmin (\edkmin)}.
\end{theorem}
\begin{proof}
Applying generalization of lemma \ref{lemmaDeltaUpper} and lemma \ref{lemmaDeltaLower} to lemma \ref{lemma2} concludes the proof.
Due to lack of space we omitted part of the generalizations, which will be given in the full paper.
\end{proof}
\qed

\bibliography{bib}

\newpage
\appendix
\section{Appendix}
\begin{lemma}
\label{lemma4}
For $i\ge 0$, $d>0$, $\epsilon \in (0,1)$, $k> d - 1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2}$ and $l\ge 2d+2$
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_{-i}\right] \le \frac{1}{i^{d+1}} \nonumber
\end{equation}

\end{lemma}
\begin{proof}
For block $b_{-i}$, $X = \{x_1,\dots,x_{m}\} $ we define $Z_j$ to be indicator variable s.t.
\begin{equation}
Z_j =
\left\{
	\begin{array}{ll}
		1 \ \ \ &  h(x_j) < (1-\epsilon i) \frac{t|U|}{m}\\
		0 \ \ \ & otherwise
	\end{array}
\right. \nonumber
\end{equation}
In addition we define $Z = \sum_jZ_j$, and $E_{-i}$ to be the expected value of $Z$. Notice that since $Z$ is sum of indicator variables $E_{-i}= (1-\epsilon i) \frac{t}{m}m = (1-\epsilon i) t$. \\
We use the above definitions to show that

\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_{-i}\right] \le \nonumber
\end{equation}
\begin{equation}
\Pr_l[ \text{number of hash values smaller than the upper boundary of block } b_{-i} \ge t] = \nonumber
\end{equation}
\begin{equation}
\Pr_l[Z \ge t] = \Pr_l[Z-E_{-i} \ge t-E_{-i}] \le \Pr_l[|Z-E_{-i}| \ge t \epsilon i)] \nonumber
\end{equation}
Using Markov inequality, assuming l is even:
\begin{equation}
\Pr_l[|Z-E_{-i}| \ge t\epsilon i] \le \frac{E(|Z-E_{-i}|^l)}{[t \epsilon i ]^l} \nonumber
\end{equation}

We use the following from lemma \ref{lemmaIndyk}:
\begin{equation}
E(|Z-E_i|^l) \le 8(6l)^{\frac{l+1}{2}}(E_i)^{\frac{l}{2}}\nonumber
\end{equation}
Thus,
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_{-i}\right] \le
 \frac{8(6l)^{\frac{l+1}{2}}(t(1-\epsilon i) )^{\frac{l}{2}}}{[t\epsilon i]^l} \nonumber
\end{equation}
In order to have
\begin{equation}
\Pr_l\left[ RANK_{t}(h(X)) \in b_{-i}\right] \le \frac{1}{i^{d+1}} \nonumber
\end{equation}
We need
\begin{equation}
\frac{8(6l)^{\frac{l+1}{2}}(t(1-\epsilon i) )^{\frac{l}{2}}}{[t\epsilon i]^l} \le \frac{1}{i^{d+1}} \nonumber
\end{equation}
or
\begin{equation}
\frac{8(6l)^{\frac{l+1}{2}}(1-\epsilon i)^{\frac{l}{2}}}{\epsilon^l i^{(l-d-1)}} \le t^{\frac{l}{2}} \nonumber
\end{equation}
\begin{equation}
\frac{8^{\frac{2}{l}}(6l)^{1+\frac{1}{l}}(1-\epsilon i)}{\epsilon^2 i^{\frac{2(l-d-1)}{l}}} \le t \nonumber
\end{equation}
choosing $l\ge 2d+2$ and substituting variables, we now need to show
\begin{equation}
\frac{8^{\frac{2}{l}}(6l)^{1+\frac{1}{l}}(1-\epsilon i)}{\epsilon^2 i} + d - 1\le k \nonumber
\end{equation}
\begin{equation}
8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2 i} - 8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}\epsilon}{\epsilon^2}+ d - 1\le k \nonumber
\end{equation}
which is satisfied for $k> d - 1 + \cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2}$
\qed
\end{proof}

\begin{lemma}
\label{lemmaDeltaLower}
For $d=2$, $\epsilon \in (0,1)$, $k>d-1 + 2\cdot8^{\frac{2}{l}}\frac{(6l)^{1+\frac{1}{l}}}{\epsilon^2}$ and $l\ge 2d+2$:\\
\begin{center}
$-\Delta \le \epsilon\frac{{k \choose d}}{{n \choose d}}$
\end{center}
\end{lemma}
\begin{proof}

\begin{equation}
\sum_{i=-\infty}^{\infty} \Pr_{l+d}\left[ RANK_{k-1}(h(X)) \in b_i\right] \times \nonumber
\end{equation}
\begin{equation}
\left[ \Pr_{l+d}\left[ h(y_1), h(y_2) \le RANK_{k-1}(h(X)) \ | \ RANK_{k-1}(h(X)) \in b_i  \right] - \frac{k}{n}\frac{k-1}{n-1} \right] \ge \nonumber
\end{equation}

We now use the lower part of the block to bound probability. 
Using $2$ independent (out of $l+d$) for $h(y_1), h(y_2),\dots,h(y_d)$, we get

\begin{equation}
\sum_{i=-\infty}^{\infty} \Pr_l\left[ RANK_{k-1}(h(X)) \in b_{i+1}\right] \left[ (\frac{k-1}{n-2})^2(1+\epsilon i)^2-\frac{k}{n}\frac{k-1}{n-1} \right] = \nonumber
\end{equation}

\begin{equation}
\sum_{i=-\infty}^{-2} \Pr_l\left[ RANK_{k-1}(h(X)) \in b_{i+1}\right] \left[ (\frac{k-1}{n-2})^2(1+\epsilon i)^2-\frac{k}{n}\frac{k-1}{n-1} \right] + \nonumber
\end{equation}
\begin{equation}
\Pr_l\left[ RANK_{k-1}(h(X)) \in b_0\right] \left[ (\frac{k-1}{n-2})^2(1-\epsilon)^2-\frac{k}{n}\frac{k-1}{n-1} \right] + \nonumber
\end{equation}
\begin{equation}
\Pr_l\left[ RANK_{k-1}(h(X)) \in b_1\right] \left[ (\frac{k-1}{n-2})^2-\frac{k}{n}\frac{k-1}{n-1} \right] + \nonumber
\end{equation}
\begin{equation}
\sum_{i=1}^{\infty} \Pr_l\left[ RANK_{k-1}(h(X)) \in b_{i+1}\right] \left[ (\frac{k-1}{n-2})^2(1+\epsilon i)^2-\frac{k}{n}\frac{k-1}{n-1} \right] = \nonumber
\end{equation}

By changing the order we get a telescoping sum as follows:
\begin{equation}
\sum_{i=-\infty}^{-2} \Pr_l\left[ RANK_{k-1}(h(X)) \in b_{\le i+1}\right] \left[ (\frac{k-1}{n-2})^2(1+\epsilon i)^2- (\frac{k-1}{n-2})^2(1+\epsilon (i+1))^2\right] + \nonumber
\end{equation}
\begin{equation}
\Pr_l\left[ RANK_{k-1}(h(X)) \in b_{\le0}\right] \left[ (\frac{k-1}{n-2})^2(1-\epsilon)^2-\frac{k}{n}\frac{k-1}{n-1} \right] + \nonumber
\end{equation}
\begin{equation}
\Pr_l\left[ RANK_{k-1}(h(X)) \in b_{\ge1}\right] \left[ (\frac{k-1}{n-2})^2-\frac{k}{n}\frac{k-1}{n-1} \right] + \nonumber
\end{equation}
\begin{equation}
\sum_{i=1}^{\infty} \Pr_l\left[ RANK_{k-1}(h(X)) \in b_{\ge i+1}\right] \left[ (\frac{k-1}{n-2})^2(1+\epsilon i)^2- (\frac{k-1}{n-2})^2(1+\epsilon (i-1))^2\right] \ge \nonumber
\end{equation}

applying lemma \ref{lemma3} and lemma \ref{lemma4}
\begin{equation}
-\sum_{i=-\infty}^{-2} \frac{1}{|i+1|^3} |(\frac{k-1}{n-2})^2(1+\epsilon i)^2- (\frac{k-1}{n-2})^2(1+\epsilon (i+1))^2| - \nonumber
\end{equation}
\begin{equation}
|(\frac{k-1}{n-2})^2(1-\epsilon)^2-\frac{k}{n}\frac{k-1}{n-1}| - |(\frac{k-1}{n-2})^2-\frac{k}{n}\frac{k-1}{n-1}| -\nonumber
\end{equation}
\begin{equation}
\sum_{i=1}^{\infty} \frac{1}{(i+1)^3} |(\frac{k-1}{n-2})^2(1+\epsilon i)^2- (\frac{k-1}{n-2})^2(1+\epsilon (i-1))^2| = \nonumber
\end{equation}

\begin{equation}
-\sum_{i=2}^{\infty} \frac{1}{(i+1)^3} |(\frac{k-1}{n-2})^2(\epsilon^2(2i-1) - 2\epsilon))| - \nonumber
\end{equation}
\begin{equation}
|(\frac{k-1}{n-2})^2(1-\epsilon)^2-\frac{k}{n}\frac{k-1}{n-1}| - |(\frac{k-1}{n-2})^2-\frac{k}{n}\frac{k-1}{n-1}| -\nonumber
\end{equation}
\begin{equation}
\sum_{i=1}^{\infty} \frac{1}{(i+1)^3} |(\frac{k-1}{n-2})^2(\epsilon^2(2i-1) + 2\epsilon))| \ge \nonumber
\end{equation}

\begin{equation}
-\frac{1}{2}\frac{k}{n}\frac{k-1}{n-1} \sum_{i=2}^{\infty} \frac{1}{(i+1)^3} |\epsilon^2(2(i+1)-1) - 2\epsilon| - \nonumber
\end{equation}
\begin{equation}
|(\frac{k-1}{n-2})^2(1-\epsilon)^2-\frac{k}{n}\frac{k-1}{n-1}| - |(\frac{k-1}{n-2})^2-\frac{k}{n}\frac{k-1}{n-1}| -\nonumber
\end{equation}
\begin{equation}
\frac{1}{2}\frac{k}{n}\frac{k-1}{n-1} \sum_{i=2}^{\infty} \frac{1}{(i+1)^3} |\epsilon^2(2(i+1)-1) + 2\epsilon| \ge \nonumber
\end{equation}

\begin{equation}
-c\frac{k}{n}\frac{k-1}{n-1}\epsilon  \nonumber
\end{equation}
\end{proof}
\qed

\begin{lemma}
\label{lemmaIndyk}
Let $Z_j$ be a set of indicator variables, let $Z = \sum_jZ_j$, let $E_i$ be the expected value of $Z$, and let $l>0$ be even. 
\begin{equation}
E(|Z-E_i|^l) \le 8(6l)^{\frac{l+1}{2}}(E_i)^{\frac{l}{2}}\nonumber
\end{equation}
\end{lemma}
\begin{proof}
The proof is based on Indyk's lemma 2.2 in \cite{Indyk99asmall}, with the following minor change:
\begin{equation}
E(|Z-E_i|^l) \le  2\sum_{j=1}^\infty(j^{l}\cdot2 e^{-\frac{j^2}{2E_i^2}E_i}) \le 4(3E_i)^\frac{l}{2}\sum_{j=1}^\infty(s^{l}e^{-s^2}) \nonumber
\end{equation}
\qed
\end{proof}

\end{document}